OpenStudy (anonymous):

Solve for x: −3|x + 7| = −12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution I think it's D

pooja195 (pooja195):

Not quite the -3 is correct but not the 11

pooja195 (pooja195):

x+7=−4(Possibility 2) x+7−7=−4−7(Subtract 7 from both sides) x=−11

OpenStudy (anonymous):

Ohh. I feel stupid now :P

pooja195 (pooja195):

lol xD yr not stupid :P

OpenStudy (anonymous):

I kept making the negatives positives

pooja195 (pooja195):

Its ok we're here to help :)

OpenStudy (anonymous):

:D

pooja195 (pooja195):

:D

OpenStudy (anonymous):

-3(x+7)=-12; reduce by (-3); x+7=4; x= 4-7; x=-3